# spacing=nonfrench

### v2

@@ Abstract
Current multi-agent Large Language Model frameworks rely on explicit orchestration patterns borrowed from human organizational structures: planners delegate to executors, managers coordinate workers, and hierarchical control flow governs agent interactions. These approaches suffer from coordination overhead that scales poorly with agent count and task complexity. We propose a fundamentally different paradigm inspired by natural coordination mechanisms: agents operate locally on a shared artifact, guided only by pressure gradients derived from measurable quality signals, with temporal decay preventing premature convergence. We formalize this as optimization over a pressure landscape and prove convergence guarantees under mild conditions.
Empirically, on meeting room scheduling across 1350 trials, pressure-field coordination outperforms all baselines: 48.5% aggregate solve rate across all difficulty levels (86.7% on easy problems) versus 12.6% for conversation-based coordination, 1.5% for hierarchical control, and 0.4% for sequential and random baselines (all pairwise comparisons $p<0.001$). Temporal decay is essential: disabling it reduces solve rate by 10 percentage points. On easy problems, pressure-field achieves 86.7% solve rate compared to 37.8% for the next-best baseline. The approach maintains consistent performance from 1 to 4 agents. Implicit coordination through shared pressure gradients outperforms explicit hierarchical control. Foundation models enable this approach: their broad pretraining and zero-shot reasoning allow quality-improving patches from local pressure signals alone, without domain-specific coordination protocols. This suggests that constraint-driven emergence offers a simpler and more effective foundation for multi-agent AI.
Keywords: multi-agent systems, emergent coordination, decentralized optimization, LLM agents
@@ 1  Introduction
Multi-agent systems built on large language models address complex task automation [15, 8, 10]. The dominant paradigm treats agents as organizational units: planners decompose tasks, managers delegate subtasks, and workers execute instructions under hierarchical supervision. This coordination overhead scales poorly with agent count and task complexity.
We demonstrate that implicit coordination through shared state outperforms explicit hierarchical control—without coordinators, planners, or message passing. Across 1350 trials on meeting room scheduling, pressure-field coordination achieves 48.5% aggregate solve rate compared to 1.5% for hierarchical control ($p<0.001$, Cohen’s $h=1.97$ on easy problems). Conversation-based coordination (AutoGen-style) achieves 12.6%, while sequential and random baselines achieve only 0.4%.
@@
    We introduce temporal decay as a mechanism for preventing premature convergence. Disabling decay reduces solve rate by 10 percentage points (from 96.7% to 86.7% in ablation studies), trapping agents in local minima.
@@
    We provide empirical evidence across 1350 trials showing: (a) pressure-field outperforms hierarchical control (48.5% vs 1.5%), (b) all comparisons with baselines are highly significant ($p<0.001$).
@@ 2.1.1 Organizational Paradigms and Dependency Management
Pressure-field coordination achieves role-free coordination: any agent can address any high-pressure region without negotiating access rights or awaiting task assignment. This contrasts sharply with traditional organizational paradigms. Horling and Lesser [9] surveyed nine such paradigms—from rigid hierarchies to flexible markets—finding that all assign explicit roles constraining agent behavior. While role assignment reduces coordination complexity by pre-structuring interactions, it introduces brittleness: role changes require protocol modifications, and role failure can cascade through the system.
Our approach instantiates Malone and Crowston’s [11] coordination framework with a critical difference: the artifact itself is the shared resource, and pressure gradients serve as dependency signals. Rather than assigning roles to manage resource access, agents share read access to the entire artifact and propose changes to high-pressure regions. Coordination emerges from pressure alignment—agents reduce local pressure, which reduces global pressure through the artifact’s shared state.
@@ 2.1.2 Distributed Problem Solving and Communication Overhead
Pressure-field coordination achieves $O(1)$ inter-agent communication overhead—agents exchange no messages. Coordination occurs entirely through shared artifact reads and writes, eliminating the message-passing bottleneck. This contrasts with the Generalized Partial Global Planning framework [3], which reduces communication from $O(n^{2})$ pairwise negotiation to $O(n\log n)$ hierarchical aggregation through summary information exchange. While Generalized Partial Global Planning represents significant progress, its explicit messages—task announcements, commitment exchanges, schedule updates—still introduce latency and failure points at scale.
@@ 2.1.3 Shared Intentions and Alignment Costs
This contrasts with the SharedPlans framework [7], which formalizes joint activity through shared mental attitudes: mutual beliefs about goals, commitments, and action sequences. The framework captures human-like collaboration but requires significant cognitive machinery—intention recognition, commitment protocols, belief revision—all computationally expensive operations that scale poorly with agent count.
Our experiments validate this analysis: pressure-field coordination eliminates the overhead of explicit dialogue by coordinating through shared artifact state. The coordination overhead of belief negotiation in explicit dialogue systems can exceed its organizational benefit for constraint satisfaction tasks. The trade-off is transparency: SharedPlans supports dialogue about why agents act; pressure-field agents react to gradients without explaining reasoning.
@@ 2.1.4 Self-Organization and Emergent Coordination
Pressure-field coordination satisfies De Wolf and Holvoet’s [2] self-organization criteria: absence of external control, local interactions producing global patterns, and dynamic adaptation. They explicitly cite “gradient fields” as a self-organization design pattern—our approach instantiates this pattern with formal guarantees.
@@
The theoretical contribution formalizes this intuition through potential game theory. Theorem 1 establishes convergence guarantees for aligned pressure systems; the Basin Separation result (Theorem 3) explains why decay is necessary to escape suboptimal basins.
@@ 2.1.5 Foundation Model Enablement
Foundation Models enable stigmergic coordination through three capabilities: (1) broad pretraining allows patch proposals across diverse artifact types without domain-specific fine-tuning; (2) instruction-following allows operation from pressure signals alone, without complex action representations; (3) zero-shot reasoning interprets constraint violations without explicit protocol training. These properties make Foundation Models suitable for stigmergic coordination—they require only local context and quality signals to generate productive actions, matching pressure-field’s locality constraints.
@@ 2.2  Multi-Agent LLM Systems
Recent work has explored multi-agent architectures for Large Language Model-based task solving. AutoGen [15] introduces a conversation-based framework where customizable agents interact through message passing, with support for human-in-the-loop workflows. MetaGPT [8] encodes Standardized Operating Procedures into agent workflows, assigning specialized roles (architect, engineer, QA) in an assembly-line paradigm. CAMEL [10] proposes role-playing between AI assistant and AI user agents, using inception prompting to guide autonomous cooperation. CrewAI [1] similarly defines agents with roles, goals, and backstories that collaborate on complex tasks.
@@ 2.3  Swarm Intelligence and Stigmergy
The concept of stigmergy—indirect coordination through environment modification—was introduced by Grassé [6] to explain termite nest-building behavior. Termites deposit pheromone-infused material that attracts further deposits, leading to emergent construction without central planning. This directly instantiates Malone and Crowston’s [11] shared resource coordination: pheromone trails encode dependency information about solution quality. Complex structures arise from simple local rules without any agent having global knowledge.
Dorigo and colleagues [5, 4] formalized this insight into Ant Colony Optimization, where artificial pheromone trails guide search through solution spaces. Key mechanisms include positive feedback (reinforcing good paths), negative feedback (pheromone evaporation), and purely local decision-making. Ant Colony Optimization has achieved strong results on combinatorial optimization problems including Traveling Salesman Problem, vehicle routing, and scheduling.
@@ 2.4  Decentralized Optimization
Potential games, introduced by Monderer and Shapley [12], are games where individual incentives align with a global potential function. A key property is that any sequence of unilateral improvements converges to a Nash equilibrium—greedy local play achieves global coordination. This provides the theoretical foundation for our convergence guarantees: under pressure alignment, the artifact pressure serves as a potential function.
Distributed gradient descent methods [13, 16] address optimization when data or computation is distributed across nodes. The standard approach combines local gradient steps with consensus averaging. While these methods achieve convergence rates matching centralized alternatives, they typically require communication protocols and synchronization. Our approach avoids explicit communication entirely: agents coordinate only through the shared artifact, achieving $O(1)$ coordination overhead.
The connection between multi-agent learning and game theory has been extensively studied [14]. Our contribution is applying these insights to Large Language Model-based artifact refinement, where the “game” is defined by pressure functions over quality signals rather than explicit reward structures.
@@ 4.1  Pressure Alignment
Definition 2 (Pressure Alignment).
@@
Definition 3 (Bounded Coupling).
@@ 4.2  Connection to Potential Games
Note that this convergence result assumes finite action spaces. In practice, patches are drawn from a finite set of Large Language Model-generated proposals per region, satisfying this requirement. For infinite content spaces, convergence to approximate equilibria can be established under Lipschitz continuity conditions on pressure functions.
@@ 5.2  Basin Quality
Theorem 2 (Basin Quality).
@@
Theorem 3 (Basin Separation).
@@ 5.3  Scaling Properties
Theorem 4 (Linear Scaling).
@@
Theorem 5 (Parallel Convergence).
Under the same alignment conditions as Theorem 1, with $K$ patches validated in parallel per tick where patches affect disjoint regions, the system reaches a stable basin within:
@@ 6  Experiments
Key findings: Pressure-field coordination outperforms all baselines (§6.2). Temporal decay is critical—disabling it reduces solve rate by 10 percentage points (§6.3). The approach maintains consistent performance from 1 to 4 agents (§6.4).
@@ 6.1.1 Task: Meeting Room Scheduling
Table 2: Problem configurations. Pre-scheduled percentage indicates meetings already placed; remaining must be scheduled by agents.

### v3

# Emergent Coordination in Multi-Agent Systems via Pressure Fields and Temporal Decay

@@ Abstract
Current multi-agent large language model (LLM) frameworks rely on explicit orchestration patterns borrowed from human organizational structures: planners delegate to executors, managers coordinate workers, and hierarchical control flow governs agent interactions. These approaches suffer from coordination overhead that scales poorly with agent count and task complexity. We propose a fundamentally different paradigm inspired by natural coordination mechanisms: agents operate locally on a shared artifact, guided only by pressure gradients derived from measurable quality signals, with temporal decay preventing premature convergence. We formalize this as optimization over a pressure landscape and prove convergence guarantees under mild conditions.
Empirically, on meeting room scheduling across 1350 total trials (270 per strategy), pressure-field coordination achieves $4\times$ higher solve rates than conversation-based coordination and over $30\times$ higher than hierarchical control (48.5% vs 11.1% vs 1.5%; all pairwise comparisons $p<0.001$). Ablation studies suggest temporal decay is beneficial. On easy problems, pressure-field achieves 86.7% solve rate compared to 33.3% for the next-best baseline. Foundation models enable this approach: their broad pretraining and zero-shot reasoning allow quality-improving patches from local pressure signals alone, without domain-specific coordination protocols. This suggests that constraint-driven emergence offers a simpler and more effective foundation for multi-agent AI.
keywords:
multi-agent systems, emergent coordination, decentralized optimization, LLM agents, stigmergy
@@ 1 Introduction
Multi-agent systems built on Large Language Models address complex task automation [wu2023autogen, hong2023metagpt, li2023camel]. The dominant paradigm treats agents as organizational units: planners decompose tasks, managers delegate subtasks, and workers execute instructions under hierarchical supervision. This coordination overhead scales poorly with agent count and task complexity.
[…]
@@
    We introduce temporal decay as a mechanism for preventing premature convergence. Ablation studies show a 10 percentage point improvement with decay enabled (96.7% vs 86.7%), directionally consistent with the theoretical prediction that decay helps escape local minima, though not statistically significant at $n=30$.
@@
    We provide empirical evidence across 1350 total trials (270 per strategy) showing: (a) pressure-field dramatically outperforms all baselines by an order of magnitude or more, (b) all pairwise comparisons are highly significant ($p<0.001$).
This work demonstrates that Foundation Model capabilities and Multi-Agent System coordination mechanisms are mutually enabling, not merely additive. Foundation Models solve a fundamental Multi-Agent System problem: traditional coordination requires explicit action space enumeration, but for open-ended artifact refinement the space of valid improvements is unbounded. Foundation Models’ broad pretraining provides implicit coverage of improvement strategies without domain-specific action representations—a “universal actor” capability. Conversely, Multi-Agent System coordination solves a fundamental Foundation Model problem: how to combine multiple model outputs coherently. Pressure gradients provide an objective criterion for output selection, replacing ad-hoc voting or ranking with principled quality-based filtering. This bidirectional synthesis explains why pressure-field coordination outperforms conversation-based alternatives that lack objective gradients for output combination.
@@ 2.1.1 Organizational Paradigms and Dependency Management
Pressure-field coordination achieves role-free coordination: any agent can address any high-pressure region without negotiating access rights or awaiting task assignment. This contrasts sharply with traditional organizational paradigms. Horling and Lesser [horling2004survey] surveyed nine such paradigms—from rigid hierarchies to flexible markets—finding that all assign explicit roles constraining agent behavior. Dignum [dignum2009handbook] systematizes this tradition, defining organizational models through three dimensions: structure (roles and relationships), norms (behavioral constraints), and dynamics (how organizations adapt). These dimensions require explicit specification and maintenance—designers must anticipate role interactions, encode coordination norms, and implement adaptation mechanisms.
Pressure-field coordination eliminates all three dimensions through gradient-based coordination. Roles dissolve: any agent may address any high-pressure region without negotiating access rights or awaiting task assignment. Norms become implicit: the pressure function encodes what “good” behavior means, and agents that reduce pressure are by definition norm-compliant. Dynamics emerge naturally: temporal decay continuously destabilizes the pressure landscape, forcing ongoing adaptation without explicit organizational change protocols. Where Dignum’s organizational models require designers to specify “who may do what with whom,” pressure-field coordination answers: “anyone may improve anywhere, and coordination emerges from shared perception of quality signals.”
Our approach instantiates Malone and Crowston’s [malone1994coordination] coordination framework with a critical difference: the artifact itself is the shared resource, and pressure gradients serve as dependency signals. Malone and Crowston identify “shared resource” management as a fundamental coordination pattern requiring protocols for access control, conflict resolution, and priority assignment. Pressure-field coordination implements this pattern through a different mechanism: rather than assigning roles to manage resource access, agents share read access to the entire artifact and propose changes to high-pressure regions. Selection and validation phases resolve conflicts implicitly—only pressure-reducing patches are applied, and the highest-scoring patch wins when proposals conflict. Coordination emerges from pressure alignment—agents reduce local pressure, which reduces global pressure through the artifact’s shared state.
@@ 2.1.2 Distributed Problem Solving and Communication Overhead
Pressure-field coordination achieves $O(1)$ inter-agent communication overhead—agents exchange no messages. Coordination occurs entirely through shared artifact reads and writes, eliminating the message-passing bottleneck. This contrasts with the Generalized Partial Global Planning framework [decker1995gpgp], which reduces communication from $O(n^{2})$ pairwise negotiation to $O(n\log n)$ hierarchical aggregation through summary information exchange. While Generalized Partial Global Planning represents significant progress, its explicit messages—task announcements, commitment exchanges, schedule updates—still introduce latency and failure points at scale.
@@ 2.1.3 Shared Intentions and Alignment Costs
This contrasts sharply with two foundational frameworks for joint activity. The SharedPlans framework [grosz1996sharedplans] formalizes collaboration through shared mental attitudes: mutual beliefs about goals, commitments, and action sequences. Cohen and Levesque’s [cohen1991teamwork] Joint Intentions theory provides an even more stringent requirement: team members must hold mutual beliefs about the joint goal, individual commitments to the goal, and mutual beliefs about each member’s commitment. Both frameworks capture human-like collaboration but require significant cognitive machinery—intention recognition, commitment protocols, belief revision—all computationally expensive operations that scale poorly with agent count.
Pressure-field coordination eliminates the mutual belief formation that Joint Intentions requires. Where Cohen and Levesque demand that each agent believe that all teammates are committed to the joint goal (and believe that all teammates believe this, recursively), pressure-field agents need only observe local pressure gradients. The shared artifact is the mutual belief—agents perceive the same pressure landscape without explicit belief exchange. This eliminates the infinite regress of “I believe that you believe that I believe” that makes Joint Intentions computationally expensive at scale.
Our experiments validate this analysis: pressure-field coordination eliminates the overhead of explicit dialogue by coordinating through shared artifact state. The coordination overhead of belief negotiation in explicit dialogue systems can exceed its organizational benefit for constraint satisfaction tasks. The trade-off is transparency: SharedPlans and Joint Intentions support dialogue about why agents act and what teammates are committed to; pressure-field agents react to gradients without explaining reasoning or maintaining models of teammate intentions.
@@ 2.1.4 Self-Organization and Emergent Coordination
Pressure-field coordination satisfies the self-organization criteria established by two complementary frameworks. De Wolf and Holvoet [dewolf2005engineering] characterize self-organizing systems through absence of external control, local interactions producing global patterns, and dynamic adaptation. They explicitly cite “gradient fields” as a self-organization design pattern—our approach instantiates this pattern with formal guarantees.
Serugendo et al. [serugendo2005selforganisation] provide a more fine-grained taxonomy, identifying four mechanisms through which self-organization emerges: (1) positive feedback amplifying beneficial behaviors, (2) negative feedback dampening harmful behaviors, (3) randomness enabling exploration, and (4) multiple interactions allowing local behaviors to propagate globally. Pressure-field coordination instantiates all four mechanisms:
  • Positive feedback: Successful patches are stored as few-shot examples (“positive pheromones”), increasing the probability of similar improvements in neighboring regions. This amplifies productive behaviors.
  • Negative feedback: Temporal decay continuously erodes fitness, preventing any region from becoming permanently “solved.” Inhibition further dampens over-activity in recently-patched regions.
  • Randomness: Stochastic model sampling and band escalation (exploitation $\to$ exploration) inject controlled randomness that prevents premature convergence to local optima.
  • Multiple interactions: Each tick produces $K$ parallel patch proposals across agents, with selection applying the best improvements. Multiple interactions per timestep accelerate the propagation of successful strategies.
@@
The theoretical contribution formalizes this intuition through potential game theory. Theorem 5.1 establishes convergence guarantees for aligned pressure systems; the Basin Separation result (Theorem 5.3) explains why decay is necessary to escape suboptimal basins. This connects self-organization principles to formal coordination theory: Serugendo’s four mechanisms map to our formal model—positive feedback to pheromone reinforcement, negative feedback to decay and inhibition, randomness to stochastic sampling, and multiple interactions to parallel validation.
@@ 2.1.5 Foundation Model Enablement
Foundation Models enable stigmergic coordination by providing capabilities that map directly to stigmergic requirements. This mapping explains why pressure-field coordination becomes practical with Foundation Models in ways it was not with prior agent architectures.
Broad pretraining enables domain-general patch proposal. Stigmergic coordination requires agents that can propose improvements based solely on local quality signals. Traditional agents require domain-specific action representations—enumerated moves, parameterized operators, or learned policies. Foundation Models’ pretraining on diverse corpora (code, text, structured data) allows patch generation across artifact types without fine-tuning. An Large Language Model can propose meeting schedule adjustments, code refactorings, or configuration changes through the same interface: observe local context, receive pressure feedback, generate improvement. This “universal actor” capability is what makes stigmergic coordination practical for open-ended artifact refinement.
Instruction-following replaces action space enumeration. Stigmergic agents need only local context to act—they should not require global state, explicit goals, or complex action grammars. Foundation Models’ instruction-following capabilities allow operation from natural language pressure descriptions alone. Rather than encoding “reduce scheduling conflicts” as a formal operator with preconditions and effects, we simply prompt: “This time block has 3 double-bookings. Propose a schedule change to reduce conflicts.” The Foundation Model interprets this pressure signal and generates appropriate patches without requiring designers to enumerate valid actions.
Zero-shot reasoning interprets quality signals. Stigmergic coordination requires agents to recognize quality deficiencies locally. Foundation Models can identify constraint violations, inefficiencies, and improvement opportunities from examples alone, without explicit training on domain-specific quality metrics. When presented with a schedule showing “Meeting A and Meeting B both have Alice at 2pm,” the model recognizes the conflict and proposes resolutions—not because it was trained on scheduling, but because conflict recognition transfers from pretraining.
In-context learning implements pheromone memory. Stigmergic systems reinforce successful strategies through positive feedback. Foundation Models’ in-context learning provides this mechanism naturally: successful patches become few-shot examples in subsequent prompts, increasing the probability of similar improvements. This “positive pheromone” effect requires no external memory system—the prompt itself carries reinforcement signals.
Generative flexibility enables unbounded solution spaces. Traditional stigmergic systems (ant colony optimization, particle swarm) operate over discrete, enumerated solution spaces. Foundation Models generate from effectively continuous spaces, proposing patches that no designer anticipated. This generative flexibility is essential for open-ended artifact refinement where the space of valid improvements cannot be enumerated in advance.
These five capabilities—domain-general patches, instruction-based operation, zero-shot quality recognition, in-context reinforcement, and generative flexibility—collectively enable stigmergic coordination for artifact refinement tasks that were previously intractable. The FM-MAS synthesis is not merely additive: Foundation Models solve the action enumeration problem that blocked stigmergic approaches, while stigmergic coordination solves the output combination problem that limits single-Foundation Model systems.
@@ 2.2 Multi-Agent LLM Systems
Recent work has explored multi-agent architectures for Large Language Model-based task solving. AutoGen [wu2023autogen] introduces a conversation-based framework where customizable agents interact through message passing, with support for human-in-the-loop workflows. MetaGPT [hong2023metagpt] encodes Standardized Operating Procedures into agent workflows, assigning specialized roles (architect, engineer, QA) in an assembly-line paradigm. CAMEL [li2023camel] proposes role-playing between AI assistant and AI user agents, using inception prompting to guide autonomous cooperation. CrewAI [crewai2024] similarly defines agents with roles, goals, and backstories that collaborate on complex tasks.
@@ 2.3 Swarm Intelligence and Stigmergy
The concept of stigmergy—indirect coordination through environment modification—was introduced by Grassé [grasse1959stigmergie] to explain termite nest-building behavior. Termites deposit pheromone-infused material that attracts further deposits, leading to emergent construction without central planning. This directly instantiates Malone and Crowston’s [malone1994coordination] shared resource coordination: pheromone trails encode dependency information about solution quality. Complex structures arise from simple local rules without any agent having global knowledge.
Dorigo and colleagues [dorigo1996ant, dorigo1997acs] formalized this insight into Ant Colony Optimization, where artificial pheromone trails guide search through solution spaces. Key mechanisms include positive feedback (reinforcing good paths), negative feedback (pheromone evaporation), and purely local decision-making. Ant Colony Optimization has achieved strong results on combinatorial optimization problems including Traveling Salesman Problem, vehicle routing, and scheduling.
@@ 2.4 Decentralized Optimization
Potential games, introduced by Monderer and Shapley [monderer1996potential], are games where individual incentives align with a global potential function. A key property is that any sequence of unilateral improvements converges to a Nash equilibrium—greedy local play achieves global coordination. This provides the theoretical foundation for our convergence guarantees: under pressure alignment, the artifact pressure serves as a potential function.
Distributed gradient descent methods [nedic2009distributed, yuan2016convergence] address optimization when data or computation is distributed across nodes. The standard approach combines local gradient steps with consensus averaging. While these methods achieve convergence rates matching centralized alternatives, they typically require communication protocols and synchronization. Our approach avoids explicit communication entirely: agents coordinate only through the shared artifact, achieving $O(1)$ coordination overhead.
The connection between multi-agent learning and game theory has been extensively studied [shoham2008multiagent]. Our contribution is applying these insights to Large Language Model-based artifact refinement, where the “game” is defined by pressure functions over quality signals rather than explicit reward structures.
@@ 4.1 Pressure Alignment
Definition 4.1 (Pressure Alignment).
@@
Definition 4.2 (Bounded Coupling).
@@ 4.2 Connection to Potential Games
Note that this convergence result assumes finite action spaces. In practice, patches are drawn from a finite set of Large Language Model-generated proposals per region, satisfying this requirement. More fundamentally, the validation phase (Phase 2b) implicitly discretizes the action space: only patches that reduce pressure are accepted, so the effective action set at any state is the finite set of pressure-reducing proposals generated that tick. For infinite content spaces, convergence to approximate equilibria can be established under Lipschitz continuity conditions on pressure functions.
@@ 5.2 Basin Quality
Theorem 5.2 (Basin Quality).
@@
Theorem 5.3 (Basin Separation).
@@ 5.3 Scaling Properties
Theorem 5.4 (Linear Scaling).
@@
Theorem 5.5 (Parallel Convergence).
Under the same alignment conditions as Theorem 5.1, with $K$ patches validated in parallel per tick where patches affect disjoint regions, the system reaches a stable basin within:
@@ 6 Experiments
Key findings: Pressure-field coordination outperforms all baselines (§6.2). Temporal decay shows a beneficial trend, though statistical significance requires larger samples (§6.3). The approach maintains consistent performance from 1 to 4 agents (§6.4). Despite using more tokens per trial, pressure-field achieves 12% better token efficiency per successful solve (§6.9).
@@ 6.1.1 Task: Meeting Room Scheduling
Table 2: Problem configurations. Pre-scheduled percentage indicates meetings already placed; remaining meetings must be scheduled by agents.
@@
Alignment verification: The per-region pressure computation uses only gaps, overlaps, and util_var—all strictly local to each time block. The unsched component is added to total pressure only, not per-region. This makes the per-region pressure separable: modifying region $i$ has zero effect on region $j$’s pressure for $j\neq i$, satisfying the alignment condition (Definition 2) with $\epsilon=0$. While attendee constraints could theoretically create cross-region coupling (the same person attending meetings in different time blocks), our overlap sensor counts overlaps only within each time block, eliminating this coupling source. Empirical analysis (Appendix B) confirms that all observed pressure improvements are positive, consistent with separable pressure.